OpenStudy (anonymous):

A chair company produces two types of chairs: Chair A and Chair B. The company is bound by the following constraints: • Chair A requires 1 hour of waxing time and 5 hours of painting time. • Chair B requires 4 hours of waxing time and 6 hours of painting time. • The maximum number of hours per week available for waxing and painting are 100 and 60, respectively. • The cost to the company is $0.79 per chair A and $2.00 per chair B. Total weekly costs cannot exceed $50.00.

OpenStudy (anonymous):

Let x = the number of Chair A produced in a week and y = the number of Chair B produced in a week. Write a system of three inequalities that describes these constraints.

OpenStudy (anonymous):

i got x + 5y ≤ 100 4x + 6y ≤ 60 .79x + 2y ≤ 50

OpenStudy (aripotta):

i could be wrong here, but wouldn't the first two be \[6x \le 100\]\[10y \le 60\]

OpenStudy (aripotta):

because x and y represent chair a and chair b, not waxing and painting

OpenStudy (aripotta):

the third inequality you have there is correct though

OpenStudy (anonymous):

Is this right? x + 5y ≥ 100 10x + 6y ≥ 60 2x + .79y ≤ 50

OpenStudy (aripotta):

no, i believe what i put down is right

OpenStudy (anonymous):

oh i didn't see sorry

OpenStudy (aripotta):

it's ok

OpenStudy (anonymous):

what you put isn't a part of the answers. x + 5y ≥ 100 10x + 6y ≥ 60 2x + .79y ≤ 50 x + 5y ≤ 100 4x + 6y ≤ 60 .79x + 2y ≤ 50 x + 4y ≤ 100 5x + 6y ≤ 60 .79x + 2y ≤ 50 x + 4y ≤ 100 5x + 6y ≤ 60 2x + .79y ≤ 50

OpenStudy (aripotta):

that doesn't make any sense

OpenStudy (aripotta):

i guess go with the second one, since it has the right numbers in it, at least

OpenStudy (anonymous):

alright thanks. I have one more, do you mind helping?

OpenStudy (aripotta):

i don't mind, i just hope it isn't wacky like this one lol

OpenStudy (anonymous):

p(x) = -x^4 – 2x^3 + 6x^2 + 8x q(x) = x + 8 Use a graphing tool to graph p(x) and q(x) to determine which of the following are the best estimate for the x values to the solutions to p(x) = q(x)? I got x = -2.0, x = -1.8, x = 0.8, x = 1.9

OpenStudy (aripotta):

i don't have a graphing tool. what are you using?

OpenStudy (anonymous):

yeah I have an app for it

OpenStudy (aripotta):

oh

OpenStudy (aripotta):

yea, i don't know. sorry :/